OpenStudy (anonymous):

Simplify: (4x − 6) − (5x + 1)

OpenStudy (anonymous):

a. x + 7 b .−x + 7 c. x − 7 d.−x − 7

OpenStudy (anonymous):

Yo, group like terms, guuurl.

OpenStudy (anonymous):

Distribute that negative sign, guuurl.

OpenStudy (anonymous):

that minus sign in the middle confuses me?

OpenStudy (anonymous):

mannnnnnnn

OpenStudy (anonymous):

Distribute it. -(5x+1) becomes -5x -1

OpenStudy (anonymous):

Once you distribute the -, you can just group like terms.

OpenStudy (anonymous):

would it be c?

OpenStudy (anonymous):

If you explain to me how you got that, I might answer that question.

OpenStudy (anonymous):

well idk if its right but i kinda just guessed , it was (4x-6)-5x-1 so added the 5 and 4 and dropped it which came just x and added the 6 and the 1 and kept the minus sign between the 5x-1 and i got x-7 ?

OpenStudy (anonymous):

I thought that you guessed =) (4x-6)-5x-1 It might help if you rewrite it like this 4x -5x -6 -1

OpenStudy (anonymous):

Na , i actually tried this timeeee loll but it was wrong? :/

OpenStudy (anonymous):

mhmm Like terms. So do the x terms together first 4x - 5x

OpenStudy (anonymous):

ugh (-_-) ohkay.

OpenStudy (anonymous):

is it A? because the opposite of subtraction is addition so it would be a + instead of -? if this aint right then i give up ..

OpenStudy (anonymous):

You need help with adding and subtracting with negatives =/

OpenStudy (anonymous):

omg -.- i give up ill just guess..

OpenStudy (anonymous):

Don't give up. Practice some negatives with me. Some examples. 5-3 = 2 3-5 = -2 4+1 = 5 -4 - 1 =-5

OpenStudy (anonymous):

Do these: 8-9 = -2 + 1 = -4 - 2 = 8-3 = 3-5 =

OpenStudy (anonymous):

is this right? -2+1 =-3?

OpenStudy (anonymous):

Okay, here's what I want you to do when one is negative and the other is positive. First, ignore the signs and just do a subtraction. 2-1 = 1. Your answer will either be 1 or -1. How do we tell which it is? We look at what sign the bigger number had in the problem. -2 + 1, okay so the bigger number is 2 and that is negative, so your answer will be negative. -2 + 1 = -1

OpenStudy (anonymous):

Another thing you can always do is picture a number line:  -2 is where I start. +1 would be going 1 to the right. -1 would be going 1 to the left.